Hi,
I need help with this problem: Find the base of a triangle with an area of 15 sq cm and 4 cm height.

Thank you

Answer by checkley75(3666)

AREA=1/2*B*H
15=1/2*B*4
15=2B
B=15/2
B=7.5 CM